OpenStudy (anonymous):

determine the volume of a 0.530 M KMno4 solution required to completely reast with 2.6 g of zn . my homework..

OpenStudy (anonymous):

5 Zn + 2 KMnO4 + 8 H2SO4 = 5 ZnSO4 + K2SO4 + 2 MnSO4 + 8 H2O Moles of Zn: 2.6/ 65 = .04 mol Volume of KMnO4: (.04 * 2/5) / .53 = .03 L